OpenStudy (anonymous):

I'm having a really hard time with this one... Does anyone know how to do this? Find the vertices, foci, and eccentricity of the ellipse. x^2 = 36 − 6y^2

OpenStudy (amistre64):

the e value is c/a as i recall from yesterday :)

OpenStudy (aravindg):

x= + or - 6 y=0

OpenStudy (anonymous):

So is the vertices (0,0) ?

OpenStudy (aravindg):

-6,0 or 6,0

OpenStudy (amistre64):

turn this into the normal looking ellipse equation x^2 = 36 − 6y^2 ; +6y^2 +6y^2 +6y^2 ---------------- x^2 + 6y^2 = 36 ; divide out the 36 /36 / 36 /36 ---------------- x^2 y^2 --- + --- = 1 36 6

OpenStudy (aravindg):

rate gud answer plzz

OpenStudy (anonymous):

What is the eccentricity?

OpenStudy (aravindg):

In mathematics, the eccentricity, denoted e or , is a parameter associated with every conic section. It can be thought of as a measure of how much the conic section deviates from being circular. In particular, The eccentricity of a circle is zero. The eccentricity of an ellipse which is not a circle is greater than zero but less than 1. The eccentricity of a parabola is 1. The eccentricity of a hyperbola is greater than 1.

OpenStudy (amistre64):

to determine eccentricity we have to know what this form tells us; the larger value is the major axis and is by convention called: a^2 the smaller is called b^2

OpenStudy (aravindg):

Any conic section can be defined as the locus of points whose distances are in a constant ratio to a point (the focus) and a line (the directrix). That ratio is called eccentricity, commonly denoted as "e." The eccentricity can also be defined in terms of the intersection of a plane and a double-napped cone associated with the conic section. If the cone is oriented with its axis being vertical, the eccentricity is where α is the angle between the plane and the horizontal and β is the angle between the cone and the horizontal. The linear eccentricity of a conic section, denoted c (or sometimes f or e), is the distance between its center and either of its two foci. The eccentricity can be defined as the ratio of the linear eccentricity to the semimajor axis a: that is, .

OpenStudy (amistre64):

a is the hypotenuse of a rt triangle formed from the vertex lengths and the foci

OpenStudy (anonymous):

What on earth? I'm even more confused now lol

OpenStudy (aravindg):

rate gud answer plzz

OpenStudy (anonymous):

I have done problems like this in the past I am just having a really hard time with this one for some reason... It's not that I don't know how to do them I just am not coming up with a good answer for this one. The answers I am getting to not seem right

OpenStudy (amistre64):

a^2 = c^2 + b^2 is how we determine c :) sqrt(36) = c^2 + sqrt(6) 6-sqrt(6) = c^2 c = \(\sqrt{6-\sqrt{6}}\) if i did it right

OpenStudy (amistre64):

the e value is defined by our c/a so we just plug those in

OpenStudy (anonymous):

Wouldn't C just be -6 and 6 without the square root

OpenStudy (amistre64):

you may be right :)

OpenStudy (amistre64):

im old and get confused so it helps when you correct my mistakes :)

OpenStudy (amistre64):

36 = c^2 + 6 30 = c^2 sqrt(30) = c looks much niver ;)

OpenStudy (anonymous):

It's fine! So A=0 B= -6, 6 C= -6, 6 ?

OpenStudy (anonymous):

Would that make the eccentricity 0?

OpenStudy (amistre64):

a b and c are just conventional names for parts of the ellipse; they are single numbers

OpenStudy (amistre64):

a is the length of the major axis vertex from the center; b is the length of the minor axis from the center; c is the distance from center to foci

OpenStudy (amistre64):

c and b make the legs of a right triangle and a is the hypotenuse

OpenStudy (anonymous):

I am working with ellipses though

OpenStudy (anonymous):

I'll figure it out... thanks

OpenStudy (anonymous):

ellipse foci | ((-sqrt(30), 0) | (sqrt(30), 0))~~((-5.47723, 0) | (5.47723, 0)) center | (0, 0) semimajor axis length | 6 semiminor axis length | sqrt(6)~~2.44949 area | 6 sqrt(6) pi~~46.1718 perimeter | 24 E(5/6)~~27.7302 focal parameter | sqrt(6/5)~~1.09545 eccentricity | sqrt(5/6)~~0.912871

OpenStudy (amistre64):

x^2 y^2 --- + --- = 1 ; is how an ellipse is set up; a and b can change position a^2 b^2 tho depending on major minor axis

OpenStudy (amistre64):

a is major and b is minor a^2 = b^2 + c^2 is the pythag thrm for the relationship between them 36 = 6 + c^2 30 = c^2 sqrt(30) = c .................................... e value is defined as c/a sqrt(30) ------- = sqrt(5/6) if i see it right sqrt(36)

OpenStudy (anonymous):

thanks!